• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name brads

    Log entry time 12:25:59 on November 30, 2009

    Entry number 303716

    This entry is a followup to: 303692

    keyword=Access update

    The problem with RHRS Q1 should be fixed (latched relay due to bad optocoupler, optocouplers have been replaced, fault cleared).

    Folts, et al. are working with Cryo to troubleshoot and (hopefully) fix some persistent flow instabilities. A plan is in place and will be executed immediately after lunch.

    Please leave the magnets ramped down until further notice so we don't interfere with the cryo tests.

    In other news, the Compton collimator has been installed and a problem with the RHRS DAQ has been resolved.